OpenStudy (shaeelynn):

A sequence is a function with domain the natural numbers 1,2,3,... using function notation, you can write the sequence An=A1+(n-1)d as A(n)=A1=(n-1)d and the sequence An=A1r^n-1 as a(n)=A1r^n-1. a) Is A(n)=A1+(n-1)d a linear function? Explain. If not, how can you adjust its definition so that it is a linear function? what is the slope? b) what type of function does A(n)=A1r^n-1 suggest? to what family of functions does this function belong? explain how it is related to the parent function of that family. Draw its graph.

OpenStudy (shaeelynn):

C. what type of function is suggested by the sum sequence S(n)=(A1(1-r^n)/(1-r)? by S(n)=n/2(A1+A(n))? explain each answer

OpenStudy (shaeelynn):

someone please help!!!

OpenStudy (maddielb):

.... What was the point of that

OpenStudy (soggylearnsandhelps):

s(n) = (n/2)*a(n) + a(1)/2 Linear with slope = (n/2) and intercept = a(1)/2

OpenStudy (soggylearnsandhelps):

Is this good??

OpenStudy (shaeelynn):

idk i need help with A,B, and C

OpenStudy (soggylearnsandhelps):

Give me 5 minutes?

OpenStudy (soggylearnsandhelps):

A. a(n) = (n-1)d + a(1) slope = (n-1)

OpenStudy (soggylearnsandhelps):

Since I have to go I will just answer them.

OpenStudy (soggylearnsandhelps):

B. a(n) = a(1)*r^(n-1) Exponential Because the exponent is n-1, the parent function is moved 1 unit to the right.

OpenStudy (soggylearnsandhelps):

and C. s(n) = (n/2)*a(n) + a(1)/2 Linear with slope = (n/2) and intercept = a(1)/2

OpenStudy (soggylearnsandhelps):

Is that good?

OpenStudy (shaeelynn):

yes thank you soo much!!